nse Common Stock Cash Dividends Service Revenue Land Miscellaneous Expense...

nse
Common Stock
Cash
Dividends
Service Revenue
Land
Miscellaneous Expense
Supplies
Supplies Expense
Wages Expense

Points:

Feedback

Check My Work

Explanation

none

X

Financial Statements

A business will construct its financial statements in a particular order because they are interrelated. This means that items formulated in an earlier statement feed into the subsequent statements, and changes to items on one financial statement can have compounding effects on the overall financial position of a company.

Which of the following is one reason the retained earnings statement is prepared after the income statement? selector 1Net income is part of the computation for ending retained earnings.

  • Retained earnings are computed on the income statement.
  • Net income is the beginning retained earnings amount.
  • Retained earnings is really an asset.
  • Net income is part of the computation for ending retained earnings.

Points:

Which of the following is one reason the retained earnings statement is prepared before the balance sheet? selector 1Ending retained earnings must be computed for the balance sheet.

  • Ending retained earnings must be computed for the balance sheet.
  • Retained earnings is really an asset.
  • Retained earnings feeds into revenues and expenses.
  • Revenues end up in the assets section.

Points:

Feedback

Check My Work

Explanation

none

X

GGE Enterprises Inc.

On November 1 of the current year, Rob Elliot invested $31,000.00 of his cash to form a corporation

A business organized under state or federal statutes as a separate legal entity.

, GGE Enterprises Inc., in exchange for shares of common stock. No other common stock was issued during November or December. After a very successful first month of operations, the retained earnings as of November 30 were reported at $5,000.00. After all transactions have been entered into the accounting equation for the month of December, the ending balances for selected items on December 31 follow. On that date, the financial statements were prepared. The balance sheet reported total assets of $53,900.00 and total stockholders' equity of $37,705.00.

Accounts

Common

Retained

Service

Wages

Rent

Supplies

Utilities

Miscellaneous

Cash

Supplies

Land

Payable

Stock

Earnings

Dividends

Revenue

Expense

Expense

Expense

Expense

Expense

You are in Column Cash? You are in Column Supplies$7,850.00 You are in Column Land$15,000.00 You are in Column Accounts Payable? You are in Column Common Stock? You are in Column Retained Earnings$5,000.00 You are in Column Dividends$6,500.00 You are in Column Service Revenue$26,250.00 You are in Column Wages Expense$6,475.00 You are in Column Rent Expense? You are in Column Supplies Expense$4,500.00 You are in Column Utilities Expense$1,310.00 You are in Column Miscellaneous Expense$460.00

Review the following questions. Place an ‘X’ in the box to indicate which financial statement(s) report the desired information. Enter the amount reported on the financial statement.

Balance

Income

Retained Earnings

Sheet

Statement

Statement

Amount

1. What is the amount reported for total liabilities and stockholders’ equity on December 31? X
2. What is the retained earnings amount reported on December 31?
3. How much does GGE Enterprises Inc. owe to its creditors?
4. How much cash is being held by GGE Enterprises Inc.?
5. By what amount did retained earnings increase or decrease during the period?
6. What is the amount of profit or loss during December?
7. What were the total expenses for December?
You are in Column Sheet8. How much was paid for rent?
Answer #1
Q 1 Which of the following is one reason the retained earnings statement is prepared after the income statement?
Option 4
Net income is part of the computation for ending retained earnings.
Ending Retained Earnings = Opening Retained Earnings + Net Income - Dividends
Q 2 Which of the following is one reason the retained earnings statement is prepared before the balance sheet?
Option 1
Ending retained earnings must be computed for the balance sheet.
Retained Earnings goes to Stockholders Equity section in Balance Sheet
Q 3 GGE Enterprises Inc.
Balance Sheet Income Statement Retained Earnings Stmt Amount Calculation Formula
1. What is the amount reported for total liabilities and stockholders’ equity on Dec 31? X      53,900 Same as Total Assets
2. What is the retained earnings amount reported on December 31? X X         6,705 =37705-31000 Equity - Common Stock
3. How much does GGE Enterprises Inc. owe to its creditors? X      16,195 =53900-37705 Total Assets - Equity
4. How much cash is being held by GGE Enterprises Inc.? X      31,050 =53900-7850-15000 Total Assets - Supplies - Land
5. By what amount did retained earnings increase or decrease during the period? X         6,705 Same as Retained Earning ending bal as open bal is 0
6. What is the amount of profit or loss during December? X X         8,205 =6705-5000+6500 End RE - Open RE + Dividends
7. What were the total expenses for December? X      18,045 =26250-8205 Revenue - Profit or loss
8. How much was paid for rent? X         7,070 =18045-6475-4500 Total Exp-Wages exp-Supp exp